### v1

# Which Backbone to Use: A Resource-efficient Domain Specific Comparison for Computer Vision

Abstract
In contemporary computer vision applications, particularly image classification, architectural backbones pre-trained on large datasets like ImageNet are commonly employed as feature extractors. Despite the widespread use of these pre-trained convolutional neural networks (CNNs), there remains a gap in understanding the performance of various resource-efficient backbones across diverse domains and dataset sizes. Our study systematically evaluates multiple lightweight, pre-trained CNN backbones under consistent training settings across a variety of datasets, including natural images, medical images, galaxy images, and remote sensing images. This comprehensive analysis aims to aid machine learning practitioners in selecting the most suitable backbone for their specific problem, especially in scenarios involving small datasets where fine-tuning a pre-trained network is crucial. Even though attention-based architectures are gaining popularity, we observed that they tend to perform poorly under low data finetuning tasks compared to CNNs. We also observed that some CNN architectures such as ConvNeXt, RegNet and EfficientNet performs well compared to others on a diverse set of domains consistently. Our findings provide actionable insights into the performance trade-offs and effectiveness of different backbones, facilitating informed decision-making in model selection for a broad spectrum of computer vision domains. Our code is available here: https://github.com/pranavphoenix/Backbones
Keywords Image Classification $\cdot$ Backbone $\cdot$ Fine tuning $\cdot$ Domains
Domains
Ranking Natural Texture Remote Sensing Plant Astronomy Medical Overall
Best ConvNeXt-Tiny ConvNeXt-Tiny ResNeXt-50 $32\times 4$d RegNetY-3.2GF WaveMix WaveMix ConvNeXt-Tiny
Better EfficientNetV2-S ResNeXt-50 $32\times 4$d EfficientNetV2-S ConvNeXt-Tiny ConvNeXt-Tiny EfficientNetV2-S RegNetY-3.2GF
Good RegNetY-3.2GF RegNetY-3.2GF ConvNeXt-Tiny ShuffleNetV2 $2.0\times$ DenseNet-161 RegNetY-3.2GF EfficientNetV2-S
Table 1: A summary of our observations showing the top 3 backbones for fine-tuning on multiple domains for image classification
1 Introduction
In most computer vision applications, particularly in image classification, practitioners commonly use a backbone feature extraction network paired with a classification head. This approach involves using the backbone to output features from the input image, which are then fed into a task-specific head for further computations. A prevalent practice in the field is to leverage backbones trained on large datasets, such as ImageNet-1k [1] and then fine-tune them on smaller datasets. This strategy is particularly useful due to the scarcity of large, domain-specific datasets, such as those in the medical applications.
Practitioners often resort to using off-the-shelf models from libraries like Torchvision [2], which offers a plethora of backbones with pre-trained ImageNet weights. These pre-trained backbone architectures, when fine-tuned on smaller, domain-specific datasets, consistently outperform models trained from scratch. This method not only enhances performance but also reduces the computational resources and time required for training, making it a dominant procedure in the field of computer vision.
However, it is important to note that while Torchvision lists the performance metrics (top-1 accuracy) of various backbone architectures on the ImageNet dataset, these metrics might not necessarily translate to similar performance rankings in fine-tuning datasets. For instance, a backbone with higher ImageNet performance might not guarantee superior performance on a particular fine-tuning dataset. Conversely, a backbone with weaker ImageNet performance might outperform others when fine-tuned on a dataset which belongs to a different domain. This possible discrepancy highlights the need for careful selection and evaluation of backbone models in various fine-tuning datasets belonging to different domains, underscoring the complexity and variability of transfer learning in computer vision.
Moreover, one of the major challenges for practitioners is the limitation in terms of resources. These limitations include GPU availability, training and inference time, and even model size. To address these challenges, our experiments focus on comparing only the light-weight (model size less than 100 MB) backbone architectures available on Torchvision, which are designed to be resource-efficient while maintaining high inference speed.
The lack of benchmarks that fit these backbones across multiple domains presents a significant challenge for practitioners in deciding which backbone to use for specific datasets with limited data. Additionally, the size of the fine-tuning dataset can also impact the performance of these backbone architectures. Our study also aims to investigate whether the dataset size influences the selection of backbone models and to provide insights on choosing the most suitable backbone based on the specific domain and the amount of available fine-tuning data.
We also analyse the architectures which perform well in each of the domain-specific datasets and try to understand their properties. This will help the research community to realise their fundamental limitations and design better architectures.
2 Related Works
Most of the research on architectures uses ImageNet benchmark to compare image classification performance of backbones. Pytorch Image Models [3] library benchmarks ImageNet classification performance across different backbone architectures. Visual Task Adaptation Benchmark [4] evaluated the performance of computer vision models on 19 tasks that span different domains. Brigato et al. [5] designed a benchmark for data-efficient image classification spaning various domains and showed that tuning hyper-parameters like learning rate, weight decay, and batch size results in a competitive baseline outperforming most specialized methods. Taher et al. [6] benchmarked various self-supervised pre-training methods, 14 pre-trained ImageNet models on 7 diverse medical tasks. Battle of backbones [7] benchmarks few pre-trained ImageNet backbones, including those pretrained by self-supervised learning and Stable Diffusion, across a diverse set of computer vision tasks ranging from image classification, Out-of-distribution generalisation, image retrieval and object detection. There is a lack of bench-marking on transfer learning performance of popular lightweight backbones across multiple domain-specific datasets.
3 Backbone Architectures
Since most of the practitioners use out of the shelf ImageNet pretrained models from the popular Torchvision library, we selected models from this library for our experiments. We used certain criteria to filter models keeping in mind our resource-efficiency constraint. So, we decided to choose models with number of parameters less than 30 million, which amounts to a storage space of around 100 MB. WaveMix [8] model was added to our list of backbones becasue WaveMix was shown to provide state-of-the-art (SOTA) performance in multiple image classification datasets belonging to many diverse domains such as EMNIST [9] and galaxy morpholohy datasets.
Torchvision has 20 backbone families with models of multiple sizes within each family making the total available models 115. After we apply our constraint of lightweight models (less than 30 M parameters), we have 53 models belonging to 14 backbone families. We decided to take the best model (highest ImageNet top-1 accuracy) from each of these families for our experiments. We also discarded models whose ImageNet top-1 accuracy was less than 75%.
The final list of selected 11 backbones for our experiments are shown in table 2. We added both Swin-Tiny [10] and SwinV2-Tiny [11] in our list because there was an under-representation of attention-based transformer architectures in our list and the only other attention based model in Torchvision was vision-transformer (ViT) whose smallest model has 86 million parameters. We wanted to also understand how attention-based models perform in resource-efficient fine-tuning under low data regime since there has been a recent trend amoung computer vision practitioners to use transformers for all tasks.
Architectures
# Params
ImageNet-1k
Top-1 Accuracy
ResNet-50 [12] 25.6 76.13
WaveMix [8] 27.9 75.32
ConvNeXt-Tiny [13] 28.6 82.52
Swin-Tiny [10] 28.3 81.47
SwinV2-Tiny [11] 28.4 82.07
EfficientNetV2-S [14] 21.5 84.23
DenseNet-161 [15] 28.7 77.14
MobileNetV3-Large [16] 5.5 75.27
RegNetY-3.2GF [17] 19.4 81.98
ResNeXt-50 [18] $32\times 4$d 25.0 81.20
ShuffleNetV2 [19] $2.0\times$ 7.4 76.23
Table 2: List of popular computer vision backbones we used for our experiments along with the model size and ImageNet accuracy for which model weights are available. All model weights are taken from Torchvision library except for WaveMix which was taken from Github.
Below are the architectures which was selected for our experiments.
ResNet [12]: ResNets are the most popular and successful backbone architectures currently in use today since its arrival almost a dacade ago. ResNet uses residual connections to allow for the training of very deep convolutional neural networks (CNN) by mitigating the vanishing gradient problem. We use ResNet-50 for our experiments.
WaveMix [8]: WaveMix is a token-mixing architecture that uses 2-dimentional discrete wavelet transform for spacial token-mixing and has been shown to provide SOTA performance in multiple image classification datasets. We use WaveMix-192/16 (level 3) for our experiments.
ConvNeXt [13]: ConvNeXt is a recent CNN architecture designed to improve upon traditional CNNs by incorporating elements from transformer models, resulting in enhanced performance and scalability. It uses depthwise convolutions, inverted bittleneck blocks and large kernels. We use ConvNeXt-Tiny in our experiments.
Swin Transformer [10]: Swin transformer was an improvement over conventional ViT which overcame the massive data requirements for training. It incorporated efficiency using hierachial representations, limiting the attention window and merging them stage by stage. We use Swin-Tiny and SwinV2-Tiny [11] for our experiments.
EfficientNet [14]: EfficientNet is a family of CNNS that optimize both model size and speed by utilizing a compound scaling method that uniformly scales network depth, width, and resolution. It incorporates advanced techniques such as progressive learning and a mix of regular and mobile convolutions. We use EfficientNetV2-S in our experiments.
Densenet [15]: DenseNet is a CNN architecture that connects each layer to every other layer in a feed-forward fashion, promoting feature reuse and reduction in number of parameters. This dense connectivity pattern helps alleviate the vanishing gradient problem and leads to improved training efficiency and accuracy. We use Densenet-161 in our experiments.
MobileNet [16]: MobileNetV3 is a CNN architecture designed for on-device and resource-constrained environments, which combines lightweight depthwise separable convolutions with squeeze-and-excitation modules. It is a highly efficient model with improved accuracy and reduced computational complexity. We use MobileNetV3-Large in our experiments.
RegNet [17]: RegNet is a family of CNN architectures that utilize a regular design space to systematically generate a diverse range of models, optimizing for both efficiency and performance. It focuses on simple, scalable structures with uniform depth, width, and group convolution patterns, incorporating features like bottleneck blocks. We use RegNetY-3.2GF for our experiments.
ResNeXt [18]: ResNeXt is a CNN architecture that extends the ResNet model by introducing a cardinality dimension, using grouped convolutions to aggregate multiple transformations, which improves performance and efficiency. We use ResNeXt-50 $32\times 4$d for our experiments.
ShuffleNet [19]: ShuffleNet is a lightweight CNN architecture designed for efficient computation on mobile devices. It uses channel shuffling and pointwise group convolution to optimize speed and accuracy. We use ShuffleNetV2 $2.0\times$ for our experiments.
4 Datasets
Since we wanted to check the performance of various backbones on data efficient fine-tuning, we decided to perform our experiments on datasets with atmost 100,000 training images. We choose 18 piblically available datasets belonging to 6 different domains such as natural, textures, remote sensing, plants, astronomy, and medical images with number of training images ranging from 1000 to 100,000 and number of classes ranging from 2 to 200. Details of datasets in each domain is given in table 3.
Datasets Domain Description
# Training
Images
# Testing
Images
# Classes
CIFAR-10 [20] Natural Images 50,000 10,000 10
CIFAR-100 [20] Natural Images 50,000 10,000 100
Tiny ImageNet [21] Natural Images (ImageNet subset) 100,000 10,000 200
Stanford dogs [22] Natural Images (Dog breeds) 12,000 8,580 120
Flowers-102 [23] Natural Images (Flower species) 2,040 6,149 102
CUB-200-2011 [24] Natural Images (Bird species) 5,994 5,794 200
Stanford Cars [25] Natural Images (Car models) 8,144 8,041 196
Food-101 [26] Natural Images (Food categories) 75,750 25,250 101
DTD [27] Texture Images 1,880 1,880 47
UC Merced Land Use [28] Remote Sensing Images 1,680 420 21
EuroSAT [29] Remote Sensing Images 18,900 8,100 10
PlantVillage [30] Plant Images 44,343 11,105 39
PlantCLEF [31] Plant Images 10,455 1135 20
Galaxy10 DECals [32] Astronomy Images (Galaxy Morphology) 15,962 1,774 10
BreakHis $40\times$ [33] Medical Images (Histopathology) 1,398 606 2
BreakHis $100\times$ [33] Medical Images (Histopathology) 1,458 632 2
BreakHis $200\times$ [33] Medical Images (Histopathology) 1,411 611 2
BreakHis $400\times$ [33] Medical Images (Histopathology) 1,276 553 2
Table 3: Details of the datasets used specifying the domain details, number of images in the training and testing set and the number of classes for classification
4.1 Natural
CIFAR-10 [20]: CIFAR-10 dataset is a widely-used benchmark dataset for image classification, consisting of 60,000 $32\times 32$ color images in 10 different classes, with 6,000 images per class divided into 50,000 training images and 10,000 test images.
CIFAR-100 [20]: CIFAR-100 dataset uses the same images as CIFAR-10, but images are distributed across 100 different classes, with 600 images per class divided into 50,000 training images and 10,000 test images. It provides a more challenging classification task compared to CIFAR-10 due to the larger number of classes.
Tiny ImageNet [21]: Tiny ImageNet dataset is a subset of the ImageNet dataset, consisting of 200 image classes with 500 training images and 50 test images per class, each resized to $64\times 64$ pixels. It is widely used for benchmarking image classification algorithms, particularly in low-resource scenarios.
Stanford Dogs [22]: Stanford Dogs dataset is a comprehensive dataset for fine-grained image classification, containing 20,580 images of 120 different dog breeds. It is widely used for benchmarking algorithms, particularly in distinguishing between closely related categories.
Flowers-102 [23]: The Flowers 102 dataset is a dataset for fine-grained image classification, consisting of 8,189 images of flowers categorized into 102 different species. Each class has between 40 to 258 images, and the dataset is commonly used to benchmark algorithms in classification tasks due to its diversity and challenging nature.
CUB-200-2011 [24]: Caltech-UCSD Birds-200-2011 is a comprehensive dataset for fine-grained image classification, consisting of 200 bird species with 11,788 annotated images. It is widely used for benchmarking algorithms in fine-grained visual recognition tasks due to its high level of granularity.
Stanford Cars [25]: The Stanford Cars dataset is a large-scale dataset for fine-grained image classification, consisting of 16,185 images of 196 classes of cars. It is widely used for evaluating and benchmarking computer vision algorithms in tasks involving fine-grained visual recognition and object detection.
Food-101 [26]: Food-101 dataset is a large-scale dataset for food classification, containing 101,000 images of 101 different food categories, with 750 training images and 250 test images per class. It is commonly used to benchmark image recognition algorithms in the context of food and culinary applications.
4.2 Texture
DTD [27]: Describable Textures Dataset (DTD) is a collection of 5,640 texture images categorized into 47 classes based on human-describable attributes. It is used to evaluate and benchmark algorithms in texture recognition and classification tasks.
4.3 Remote Sensing
UC Merced Land Use [28]: UC Merced Land Use dataset is a high-resolution dataset for land use classification, containing 2,100 aerial images categorized into 21 land use classes with 100 images per class. Each image is $256\times 256$ pixels, and the dataset is commonly used for evaluating and benchmarking algorithms in remote sensing and geospatial analysis tasks.
EuroSAT [29]: The EuroSAT dataset is a benchmark dataset for land use and land cover classification, consisting of 27,000 RGB and multispectral images covering 10 classes, with images derived from Sentinel-2 satellite data. It is widely used for evaluating the performance of machine learning algorithms in remote sensing and geospatial analysis tasks.
4.4 Plant
PlantVillage [30]: PlantVillage dataset is a comprehensive dataset for plant disease classification, containing over 54,000 images of healthy and diseased leaves across 39 different plant categories. It is widely used for benchmarking machine learning algorithms in agricultural and plant pathology applications.
PlantCLEF [31]: The PlantCLEF dataset is a large-scale dataset for plant identification, comprising millions of images covering thousands of plant species, including trees, flowers, fruits, and leaves. It is used for evaluating and benchmarking algorithms in botanical classification and plant biodiversity studies. We use a subset of this dataset.
4.5 Astronomy
Galaxy 10 DECals [32]: Galaxy 10 DECals dataset is a dataset for galaxy classification, consisting of 17,000 images of galaxies classified into 10 different morphological categories. It is used for evaluating and benchmarking machine learning algorithms in astronomy and astrophysical research.
4.6 Medical
BreakHis [33]: Breast Cancer Histopathological Database is a dataset specifically designed for the classification of breast cancer histopathological images. It contains 7,909 microscopic images of breast tumor tissue, divided into benign and malignant categories. It provides microscopic images of breast tumor tissue at four different magnification levels: $40\times$, $100\times$, $200\times$ and $400\times$. Each magnification level offers a different level of detail, allowing for a comprehensive analysis of histopathological features. The dataset is widely used for evaluating and benchmarking algorithms in medical image analysis and computer-aided diagnosis.
5 Experimental Details
Since the task is to measure fine-tuning performance of these backbones in image classification via transfer learning, a standard training protocol was used to compare the performance of all models. We do not freeze any layers of the pre-trained backbones and fine-tune the full model on the given dataset or a fraction of it, and measure the top-1 accuracy in the test set. Only the final linear layer of the backbone was modified to match the number of classes of each fine-tuning dataset.
All images were resized to $256\times 256$ for our experiments, except for BreakHis fine-tuning where we resized images to $672\times 448$ since reducing the resolution of histopathology images lead to poor results across models. TrivialAugment [34] was used as data augmentation for all datasets except BreakHis, Galaxy 10 DEcals, Augmentation was only used after checking whether using this augmentation improved the model performance or not.
Due to limited computational resources, we use early stopping where training is stopped if accuracy stops improving after 10 epochs. All experiments were done on a single 80 GB Nvidia A100 GPU. Training was done using DualOpt [35] where we used AdamW optimizer ($\alpha=0.001,\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$) with a weight decay of 0.01 during the initial phase and then used SGD with a learning rate of $0.001$ and momentum =0.9=0.9 during the final phase.
Cross-entropy loss was used for image classification. Batch-size was choosen so that the entire GPU will be utilized while training. We used automatic mixed precision in PyTorch during training. Top-1 accuracy on the test set of best of three runs with random initialization is reported as a generalization metric based on prevailing protocols [36].
6 Results
The results of our fine-tuning experiments on datasets of all domains are shown in Table 4, Table 5 and Table 6. We find that ConvNeXt-Tiny outperforms all the other models in almost all natural image datasets (except CIFAR-10 and CIFAR-100). EfficientNetV2-S also performs well on atleast 5 out of 7 datasets. RegNetY-3.2GF also performs well on 5 of the 7 datasets. WaveMix performs the best in CIFAR-10 and CIFAR-100 but could not replicate the same performance on other datasets.
ConvNeXt retains the good performance even amoung other domains such as textures, plant and astronomy. RegNet also performs well on texture, remote sensing and plant domains. We observe that SwinV2-Tiny performs the best in UC Merced Land Use dataset, WaveMix and ResNeXT top the EuroSat dataset, Densenet tops the PlantVillage dataset and WaveMix significantly outperforms all the other models in Galaxy10 DECals dataset.
Natural Images
Backbones Stanford Dogs Flowers102 CUB200 Stanford Cars Tiny ImageNet CIFAR10 CIFAR100
ResNet50 82.03 90.77 75.34 90.84 75.33 94.58 79.05
WaveMix 83.12 88.43 67.32 87.86 76.30 97.26 83.88
ConvNeXt-Tiny 89.47 94.41 81.92 92.26 83.42 96.48 82.60
Swin-Tiny 83.81 90.07 78.16 90.04 76.85 95.38 81.25
SwinV2-Tiny 84.44 90.06 66.89 91.38 73.93 94.52 75.37
EfficientNetV2-S 86.59 93.65 79.10 91.59 81.35 96.42 83.20
DenseNet-161 80.02 89.96 76.44 91.06 75.03 96.03 81.01
MobileNetV3-large 78.85 89.36 73.91 85.50 76.74 95.92 79.51
RegNetY-3.2GF 85.94 92.22 81.24 91.33 80.14 96.82 82.89
ResNeXt-50 $32\times 4$d 85.09 91.78 78.29 90.22 77.86 96.16 81.30
ShuffleNetV2 $2.0\times$ 78.11 91.54 74.73 88.46 76.43 96.30 83.32
Table 4: Top-1 classification accuracy for fine-tuning pretrained backbones on natural image datasets. The top 3 results for each dataset are highlighted with best in green and bold, second best in lighter green, and third best in lightest green.
Domain Texture Remote Sensing Plant Astronomy
Backbones DTD
UC Merced
Land Use
EuroSat PlantVillage PlantCLEF Galaxy10
ResNet50 68.21 96.90 98.75 99.83 80.60 84.80
WaveMix 68.25 97.72 98.96 99.79 79.52 95.10
ConvNeXt-Tiny 73.70 98.33 98.74 99.80 82.71 87.29
Swin-Tiny 70.15 97.86 98.52 99.70 79.03 84.64
SwinV2-Tiny 68.78 98.81 98.50 99.76 78.16 83.15
EfficientNetV2-S 70.18 98.22 98.88 99.81 81.09 84.75
DenseNet-161 66.14 97.08 98.83 99.88 76.49 86.70
MobileNetV3-large 68.69 97.14 98.72 99.80 75.52 82.47
RegNetY-3.2GF 71.16 98.33 98.69 99.83 82.76 82.47
ResNeXt-50 $32\times 4$d 72.73 98.33 98.96 99.75 80.32 85.82
ShuffleNetV2 $2.0\times$ 68.64 97.86 98.65 99.70 81.76 83.69
Table 5: Top-1 classification accuracy for fine-tuning pretrained backbones on different domain datasets. The top 3 results are highlighted with best in green and bold, light green, and third best in lightest green.
Medical Images (Histopathology) BreakHis Dataset
Backbones $40\times$ $100\times$ $200\times$ $400\times$ Average
ResNet50 97.91 99.53 99.22 98.44 98.78
WaveMix 99.42 99.42 99.38 99.35 99.39
ConvNeXt-Tiny 95.10 90.73 88.59 88.72 90.79
Swin-Tiny 88.11 93.89 90.00 82.61 88.65
SwinV2-Tiny 89.96 92.44 88.65 83.83 88.72
EfficientNetV2-S 99.44 99.42 99.11 99.06 99.26
DenseNet-161 98.62 99.24 99.28 98.20 98.84
MobileNetV3-Large 99.56 98.24 99.50 98.94 99.06
RegNetY-3.2GF 99.84 99.22 99.48 98.02 99.14
ResNeXt-50 $32\times 4$d 99.46 99.22 99.34 98.05 99.02
ShuffleNetV2 $2.0\times$ 99.38 98.59 99.22 98.02 98.80
Table 6: Top-1 classification accuracy for fine-tuning pretrained backbones on medical datasets. The top 3 results are highlighted with best in green and bold, light green, and third best in lightest green.
7 Observations
$\bigtriangleup$ Higher pre-training accuracy in ImageNet does not mean higher fine-tuning accuracy. From Table 2 we can see that amoung all the backbones we used for our experiments, EfficientNetV2-S has the highest pre-training ImageNet-1k accuracy (84.23%). Even though the fine-tuning performance of EfficientNet was decent, it did not perform the best in any of the domains or datasets we evaluated on. Even amoung datasets such as Tiny ImageNet and Stanford Dogs whose images are sourced from ImageNet dataset, we find that ConvNeXt outperforms EfficientNet. Therefore, we recommend the practitioners to not use the pre-training accuracy as a metric to choose the backbone.
$\bigtriangleup$ Convolutional models strongly outperform transformers in resource-efficient low-data fine-tuning. Even though Swin transformer has a hierarchical structure capable of exploiting the spacial inductive bias [7], they still perform poorly in almost all our tasks compared to modern CNN architectures such as ConvNeXt (which was designed based on Swin transformer). Hence, we recommend that for the low data fine-tune regime, it is better to avoid using transformer architectures like Swin and use pure CNN backbones such as ConvNeXt, EfficientNet or RegNet.
$\bigtriangleup$ ConvNeXt architecture consistently outperforms other models when fine-tuning on natural image datasets. This superior performance can be attributed to the innovative design of ConvNeXt, which integrates architectural advancements that bridge the gap between traditional convolutional networks, like ResNet, and modern Swin transformer models. ConvNeXt retains the beneficial convolutional inductive bias, enabling it to learn more effectively than other convolution-based architectures and even some attention-based transformer models. Our data clearly indicates that for natural images, ConvNeXt stands out as the best model, delivering exceptional fine-tuning performance.
$\bigtriangleup$ RegNet and EfficientNet models are excellent choices for fine-tuning across a wide range of image domains. While ConvNeXt excels predominantly with natural images, EfficientNet closely follows in performance, and RegNet also shows strong results in this domain. However, the versatility of RegNet and EfficientNet extends beyond natural images. Our experiments reveal that these models also perform exceptionally well on diverse datasets, including remote sensing images, plant datasets, and medical images, such as histopathology images. Therefore, we recommend practitioners to consider RegNet and EfficientNet when working with datasets beyond natural images, as their adaptability and robust performance across various domains make them valuable tools for fine-tuning tasks.
$\bigtriangleup$ ShuffleNet is a better choice than MobileNet when very light-weight models are needed. When we observe the performance of very lightweight models, specifically those with a model size of less than 50 MB, which are ideal for on-device applications, we find that ShuffleNetV2 generally outperforms MobileNetV3 across multiple domains. Although MobileNetV3 has shown better performance on medical domain, ShuffleNetV2 demonstrates more consistent and slightly superior performance across a broader range of image domains. Therefore, we recommend ShuffleNet v2 as a better choice for practitioners dealing with on-device applications, where fine-tuning a model on a domain-specific dataset is required.
$\bigtriangleup$ WavMix performs well in datasets where multi-resolution token-mixing aid in learning WaveMix outperforms all other models (9% increase from the second best, ConvNeXt) in galaxy morphology classification. WaveMix also performs well in medical domain performing better than all other datasets, and also maintaining the performance across different magnification. WaveMix, which uses 2D-DWT might possess inductive bias that can analyse the domains of astronomy and medical images better than other convolutional models due its multi-resolution token-mixing. We recomment using WaveMx in domains where features across different resolutions are needed for better performance. WaveMix also performs better than ConvNeXt in CIFAR-10 and CIFAR-100 datasets, which were actually low resolution images ($32\times 32$) which were resized to ($256\times 256$) for our training. WaveMix is also state-of-the-art in many low-resolution datasets such as EMNIST ($28\times 28$). So, we recommend using WaveMix for low-resolution image datasets.
$\bigtriangleup$ Age of ResNet dominance is over. Our results also reveal that ResNet is no longer competitive compared to these modern architectures in any domain. Our experimental data shows that for natural images, ResNet does not even rank among the top three performers. It significantly lags behind newer models such as ConvNeXt. While ResNet does perform relatively well on one medical dataset, even in these cases, other models achieve similar performance levels. Therefore, it is clear that practitioners should transition from using ResNet to these newer architectures to achieve better results in their fine-tuning tasks.
$\bigtriangleup$ ConvNeXt and Swin transformers perform poorly in medical domain Our results show that both the architectures perform perform really bad compared to all the other models when used in medical domain. Therefore, we suggest practioners working in medical domain to not use these models despite their high performance in other domains.
8 Conclusions
Computer vision backbones are critical for fine-tuning on domain specific datasets, particularly when using pre-trained backbones available in libraries such as Torchvision. In our study, we have compared the performance of various lightweight, resource-efficient backbones across different domains, including medical images, natural images, astronomy images, plant images, and remote sensing images.
Our analysis revealed that modern architectural models like ConvNeXt, EfficientNet, and RegNet excel in handling multiple domains images. Additionally, models with specific inductive biases, such as WaveMix, are particularly useful for tasks requiring multi-resolution analysis. Amoung lightweight, on-device models, ShuffleNet consistently outperformed MobileNet in fine-tuning tasks. We also observed that transformer-based or attention-based models, such as Swin Transformer, do not perform well when the fine-tuning dataset is small.
Based on our findings, we offer practical recommendations for using pre-trained computer vision backbones. Our comparative analysis on image classification fine-tuning serves as a valuable guide for practitioners and researchers alike, who aim to optimize their models for various image domains. We hope that our work will contribute to the development of better model architectures capable of performing well across diverse image datasets.
Limitations: We restricted our comparison to models available in Torchvision, focusing specifically on lightweight and resource-efficient architectures. Consequently, we did not analyze any larger models with more than 30 million parameters, limiting our ability to test the scalability of these models with larger fine-tuning datasets. Furthermore, our analysis was confined to fine-tuning datasets containing fewer than 100,000 images, which may not fully represent scenarios involving significantly larger datasets and scalability of these backbones.
Another limitation of our work is that we exclusively focused on the computational task of image classification. We did not extend our analysis to other important tasks in computer vision, such as object detection or image retrieval. The performance of various backbones on these other tasks remains unexplored in our study. While we hope that there might be some correlation with the performance we observed to these other computer vision tasks, this remains speculative and requires further investigation to confirm.
Computation Cost and Carbon Footprint: The experiments in this paper took a cumulative 1500 GPU hours on NVIDIA RTX A100 cards. Assuming the GPUs were running with an average carbon efficiency of 0.37 kgCO2eq/kWh, the total emissions are estimated to be 222 kgCO2eq.